OpenStudy (ally_b):

Determine two pairs of polar coordinates for the point (4, 4) with 0° ≤ θ < 360°. (4 square root 2, 135°), (-4 square root 2, 315°) (4 square root 2, 45°), (-4 square root 2, 225°) (4 square root 2, 315°), (-4 square root 2, 135°) (4 square root 2, 225°), (-4 square root 2, 45°)

OpenStudy (sshayer):

\[put~4=r\cos \theta~and~ 4=r \sin \theta\] square and add \[16+16=32=r^2\left( \cos ^2\theta+\sin ^2 \theta \right),r^2=32,r=\pm \sqrt{32}\] as x and y co-ordinates are positive so point lies in first co-ordinate. divide \[\frac{ r \sin \theta }{ r \cos \theta }=\frac{ 4 }{ 4 }=1,\tan \theta=1,\theta =45\] so polar co-ordinates are (r,theta) and (-r,theta +180) or \[(4\sqrt{2},45),(-4,45+180)~or~(4\sqrt{2},45) ~and~(-4,225)\]

OpenStudy (ally_b):

So the answer is B? @sshayer

OpenStudy (sshayer):

pair of polar co-ordinates are=\[\left( r,\theta \right) and \left( -r,\theta \pm \pi \right)\] second option is correct.

OpenStudy (ally_b):

Thank you

OpenStudy (sshayer):

yw